OpenStudy (anonymous):

sovle 5 7/12+2 1/6 sovle 5 7/12+2 1/6 @Mathematics

OpenStudy (anonymous):

\[\frac{93}{12} = 7\frac{9}{12} \]

OpenStudy (anonymous):

\[also, \frac{93}{12} = \frac{31}{4} = 4\frac{3}{4}\]

OpenStudy (anonymous):

\[\left(5+\frac{7}{12}\right)+\left(2+\frac{1}{6}\right)=\frac{31}{4}=7\frac{3}{4} \]